OpenStudy (anonymous):

Can someone please help me with this? Let f(x)=2x^2+x-3 and g(x)=x-1. Perform the indicated operation, then find the domain. (f-g)(x) a. x^2-4; domain: positive real numbers b. 2x^2-4; domain: all real numbers c. x^2-4; domain: all real numbers d. 2x^2; domain: all real numbers

OpenStudy (retireed):

This looks complicated but it is very logical. Unless I am missing something Just take f(x) - g(x) f(x)=2x^2+x-3 - g(x)=x-1 (f-g)(x) = 2x^2+x-3 - x-1 = 2x^2- 4 You can figure out the domain.

OpenStudy (anonymous):

The domain is all real numbers. Thank you very much that was super helpful. @retirEEd

OpenStudy (anonymous):

If you don't mind... It is now asking me to use the same equation but perform it with the equation (f g)(x) @retirEEd

OpenStudy (anonymous):

I think you would set it up like this: (f g)(x) = 2x^2+x-3 x-1 (multiplying them)

OpenStudy (mathmale):

Yes, but show multiplicaion of f and g as (fg)(x) or (f*g)(x).